OpenStudy (abcdefghijklmq):

A scale model of an airplane is 7.5 inches long. The scale used to make the model is 1 inch : 5 feet. How long is the original airplane? a. 1.5 inches c. 37.5 inches b. 5 feet d. 37.5 feet

OpenStudy (zpupster):

use a ratio 7.5in/1in = x/5ft or 7.5in/1in = x/60in or 7.5in(60in)/1in = x divide by 12 to get in feet